# Properties of squeezing functions on $h$-extendible domains

### v1

@@ Abstract.
The purpose of this article is twofold. First, we prove that the squeezing function approaches 1 near strongly pseudoconvex boundary points of bounded domains in $\mathbb{C}^{n+1}$. Second, we show that the squeezing function approaches 1 along certain sequences converging to pseudoconvex boundary points of finite type in the sense of [10], including uniformly $\Lambda$-tangential and spherically $\frac{1}{2m}$-tangential convergence patterns.
@@ 1. Introduction
In this paper, we investigate the phenomenon of the squeezing function approaching $1$ along sequences converging to pseudoconvex boundary points by the scaling technique, introduced by S. Pinchuk (cf. [36]). For a smoothly bounded planar domain $D$ and $p\in\partial D$, one always has $\lim\limits_{z\to p}\sigma_{D}(z)=1$ (see [11]). For pseudoconvex domains of higher dimension, this result holds when $D$ is a bounded strongly pseudoconvex domain in $\mathbb{C}^{n+1}$ (cf. [12]), or when $p$ is globally strongly convex in the sense of [12], or when $p$ is a spherically extreme boundary point in the sense of [23].
The first aim of this paper is to prove the following theorem, which shows that the squeezing function approaches $1$ along any sequence converging to a strongly pseudoconvex boundary point.
@@
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n+1}$ and $\xi_{0}\in\partial\Omega$. If $\partial\Omega$ is $\mathcal{C}^{2}$ strongly pseudoconvex at $\xi_{0}$, then we have $\lim\limits_{z\to\xi_{0}}\sigma_{\Omega}(z)=1$.
@@
We note that A. Zimmer [42] first proved that if $\lim_{z\to\partial D}\sigma_{D}(z)=1$ for a $\mathcal{C}^{\infty}$-smooth bounded convex domain $D$, then the domain is necessarily strictly pseudoconvex. Subsequently, in contrast to Problem 1, a bounded convex $\mathcal{C}^{2}$-smooth domain $\Omega\subset\mathbb{C}^{n+1}$ that is not strongly pseudoconvex was constructed in [17], satisfying $\lim_{z\to\partial\Omega}\sigma_{\Omega}(z)=1$. Moreover, the authors [33, Example $1.2$] showed that $\sigma_{E_{1,2}}(\eta_{j})\to 1$ as $j\to\infty$ for the sequence
@@
converging tangentially to the non-strictly pseudoconvex boundary point $(0,1)$, where $E_{1,2}:=\{(z_{1},z_{2})\in\mathbb{C}^{2}\colon|z_{2}|^{2}+|z_{1}|^{4}<1\}$ (see also [33, Theorem $1.10$] for the general case). In what follows, we demonstrate that this phenomenon occurs for much larger classes of domains.
The second aim of this paper is to prove the following theorem, which tells us that the squeezing function approaches $1$ along a sequence converging uniformly $\Lambda$-tangentially to a strongly $h$-extendible boundary point (cf. Definition 4.2 and Definition 4.1, respectively).
@@
Let $\Omega$ be a bounded pseudoconvex domain in $\mathbb{C}^{n+1}$ with $C^{\infty}$-smooth boundary. Let $\xi_{0}\in\partial\Omega$ be strongly $h$-extendible with Catlin’s finite multitype $(2m_{1},$ $\ldots,2m_{n},1)$ and let $\Lambda=(1/2m_{1},\ldots,1/2m_{n})$ (see Definition 4.2). If $\{\eta_{j}\}\subset\Omega$ is a sequence converging uniformly $\Lambda$-tangentially to $\xi_{0}\in\partial\Omega$ (see Definition 4.1), then $\sigma_{\Omega}(\eta_{j})\to 1$ as $j\to\infty$.
@@
Now we turn to bounded pseudoconvex domains in $\mathbb{C}^{2}$. First of all, let us consider the following problem, first posed in [17, Problem $4.3$].
@@
Let $\Omega$ be a bounded domain in $\mathbb{C}^{2}$ and $\xi_{0}\in\partial\Omega$. Suppose that $\partial\Omega$ is $\mathcal{C}^{\infty}$-smooth, pseudoconvex and of D’Angelo finite type near $\xi_{0}$. If $\{\eta_{j}\}\subset\Omega$ is a sequence converging spherically $\frac{1}{2m}$-tangentially to $\xi_{0}\in\partial\Omega$, then $\sigma_{\Omega}(\eta_{j})\to 1$ as $j\to\infty$, where $2m$ is the type of $\partial\Omega$ at $\xi_{0}$.
@@ 3.1. Geometry of strongly pseudoconvex hypersurfaces
Therefore, passing to a subsequence if necessary, we may assume that the sequences $\Omega_{j}:=\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(\Omega)$ and $\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(U_{0}^{-})$ converge normally to the Siegel half-space
@@
Next, let us consider the sequence of biholomorphic map $f_{j}:=\Psi\circ\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}\colon\Omega\to f_{j}(\Omega)=\Psi(\Omega_{j})$. Since $\Psi(0^{\prime},-1)=(0^{\prime},0)$ and $\Psi(z,w)\to(0^{\prime},-1)$ as $\mathcal{U}_{n+1}\ni(z,w)\to\infty$, it follows that for a sufficiently small $\epsilon>0$, there exists $j_{0}\in\mathbb{N}_{\geq 1}$ such that
$f_{j}(\Omega\setminus U_{0})\subset B((0^{\prime},-1),\epsilon/2)\quad\text{for all }j\geq j_{0}.$
In addition, one notices that $f_{j}(\Omega\cap U_{0})$ converges normally to $\mathbb{B}^{n+1}$ and $f_{j}(\partial\Omega\cap U_{0})$ converges to $\partial\mathbb{B}^{n+1}$. Moreover, since $f_{j}(\eta_{j})=\Psi(0^{\prime},-1)=(0^{\prime},0)$ we may assume that
$B\big{(}(0^{\prime},0),1-\epsilon\big{)}\subset f_{j}(\Omega)\subset B\big{(}(0^{\prime},0),1+\epsilon\big{)},\quad\forall j\geq j_{0}.$
This yields that
$\sigma_{\Omega}(\eta_{j})\geq\frac{1-\epsilon}{1+\epsilon},\quad\forall j\geq j_{0},$
Since $\epsilon>0$ is arbitrary, we conclude that $\lim_{j\to\infty}\sigma_{\Omega}(\eta_{j})=1$, and thus the proof of Theorem 1.1 is complete. $\Box$
@@ 4.1. $\Lambda$-tangential convergence
and then we consider the sequence $\{Q_{j}\}$ of automorphisms of $\mathbb{C}^{n+1}$, given by
@@
Therefore, the sequence of domains $T_{j}(U_{0}^{-})$ converges normally to the following model
$M_{H}:=\left\{(\tilde{z},\tilde{w})\in\mathbb{C}^{n+1}\colon\hat{\rho}(\tilde{z},\tilde{w}):=\mathrm{Re}(\tilde{w})+H(\tilde{z})<0\right\}.$
Furthermore, one observes that $\Omega_{j}:=T_{j}(\Omega)$ converges also normally to $M_{H}$.
One notes that $M_{H}$ is also the limit of a sequence of the pseudoconvex domains $T_{j}(U_{0}^{-})$. Hence, $M_{H}$ is also pseudoconvex, and thus $H$ is plurisubharmonic. In addition, it follows directly from Lemma 4.1 that $H$ is positive definite. Therefore, there exits a biholomorphism $\Theta\colon M_{H}\to\mathcal{U}_{n+1}$, where $\mathcal{U}_{n+1}$ is the Siegel half-space, given by
@@
Now let us consider the sequence of biholomorphic map $f_{j}:=\Psi\circ\Theta\circ\Delta_{j}\circ Q_{j}\circ L_{\eta_{j}^{\prime}}\colon\Omega\to f_{j}(\Omega)=\Psi\circ\Theta(\Omega_{j})$. Since $\Theta(0^{\prime},-1)=(0^{\prime},-1),\Psi(0^{\prime},-1)=(0^{\prime},0)$, and $\Psi(z,w)\to(0^{\prime},-1)$ as $\mathcal{U}_{n+1}\ni(z,w)\to\infty$, it follows that for a sufficiently small $\epsilon>0$, there exists $j_{0}\in\mathbb{N}_{\geq 1}$ such that
$f_{j}(\Omega\setminus U_{0})\subset B((0^{\prime},-1),\epsilon/2)\quad\text{for all }j\geq j_{0}.$
Furthermore, one observes that $f_{j}(\Omega\cap U_{0})$ converges normally to $\mathbb{B}^{n+1}$ and $f_{j}(\partial\Omega\cap U_{0})$ converges to $\partial\mathbb{B}^{n+1}$. Moreover,
@@
Therefore, we may assume that
$B\big{(}(0^{\prime},0),1-\epsilon\big{)}\subset F_{j}(\Omega)\subset B\big{(}(0^{\prime},0),1+\epsilon\big{)},\quad\forall j\geq j_{0},$
where $F_{j}(.):=f_{j}(.)-f_{j}(\eta_{j}),\;\forall j\geq j_{0}$. Since $F_{j}(\eta_{j})=0$, it follows that
$\sigma_{\Omega}(\eta_{j})\geq\frac{1-\epsilon}{1+\epsilon},\quad\forall j\geq j_{0}.$
Since $\epsilon>0$ is arbitrary, we conclude that $\lim_{j\to\infty}\sigma_{\Omega}(\eta_{j})=1$, and thus the proof of Theorem 1.2 is now complete. $\Box$
@@
$\rho(z_{1},z_{2},w)=\mathrm{Re}(w)+|z_{1}|^{4}+|z_{1}|^{2}|z_{2}|^{4}+|z_{2}|^{8}+R(z_{1},z_{2}),$
where $R(z_{1},z_{2})$ is a $\mathcal{C}^{\infty}$-smooth and $R\in\mathcal{O}(\Lambda,1)$ with $\Lambda=(\frac{1}{4},\frac{1}{8})$. Then we have
$\liminf_{j\to\infty}\sigma_{\Omega}(\eta_{j})>0,$
where $\eta_{j}=\big{(}\frac{1}{j^{1/4}},\frac{1}{j^{3/8}},-\frac{1}{j}-\frac{2}{j^{2}}-\frac{1}{j^{3}}\big{)}\in\Omega,\;\forall j\in\mathbb{N}_{\geq 1}$.
@@
This yields that the sequence of domains $\Omega_{j}:=\phi_{{\eta}_{j}}(\Omega)$ converges normally to the following model
@@
Finally, let us consider the sequence of biholomorphic map $f_{j}:=\Psi\circ\Theta\circ\Phi_{\eta_{j}}\colon\Omega\to f_{j}(\Omega)$. Since $\Psi\circ\Theta\circ\phi_{{\eta}_{j}}(\eta_{j}^{\prime})=\Psi(0,1,-1)=(0,1,0)\in\partial\mathcal{D}_{1,1,2}$ and $\Psi\circ\Theta\circ\phi_{{\eta}_{j}}(\eta_{j})=\Psi(0,1,-2)=\Big{(}0,\sqrt{\dfrac{2}{3}},-\dfrac{1}{2}\Big{)}\in\mathcal{D}_{1,1,2}$. Furthermore, we have $\Psi(z_{1},z_{2},w)\to(0,0,-1)$ as $\mathcal{E}_{1,1,2}\ni(z_{1},z_{2},w)\to\infty$. We note that $f_{j}(\Omega)$ converges normally to $\mathcal{D}_{1,1,2}$ and $f_{j}(\partial\Omega)$ converges to $\partial\mathcal{D}_{1,1,2}$. Therefore, this implies that
$\sigma_{\Omega}(\eta_{j})=\sigma_{f_{j}(\Omega)}\Big{(}0,\sqrt{2/3},-1/2\Big{)}>\dfrac{1}{2}\dfrac{\mathrm{dist}\Big{(}\big{(}0,\sqrt{2/3},-1/2\big{)},\partial\mathcal{D}_{1,1,2}\Big{)}}{\mathrm{diam}(\mathcal{D}_{1,1,2})}>0$
for any $j$ big enough. Hence, the proof is now complete.
@@
Consider the domain $E_{1,2,4}$ and the sequence $\{\eta_{j}\}\subset E_{1,2,4}$ as in Proposition 4.2. Then we have $\sigma_{E_{1,2,4}}(\eta_{j})\not\to 1$ as $j\to\infty$, contrary to Theorem 1.2. Indeed, suppose, for the sake of contradiction, that $\sigma_{E_{1,2,4}}(\eta_{j})\to 1$ as $j\to\infty$, then by the argument as in the proof of [31, Theorem $2.1$] the unit ball $\mathbb{B}^{n+1}$ is biholomorphically equivalent to $\mathcal{D}_{1,1,2}=\left\{(z_{1},z_{2},w)\in\mathbb{C}^{3}\colon|w|^{2}+|z_{1}|^{2}+|z_{2}|^{4}<1\right\}$. Therefore, we arrive at a contradiction, as $\mathcal{D}_{1,1,2}$ is not homogeneous.

### v2

@@ Abstract.
The purpose of this article is twofold. We first prove that the localized squeezing function approaches 1 near strongly pseudoconvex boundary points of bounded domains in $\mathbb{C}^{n+1}$. Then, we show that the localized squeezing function approaches 1 along certain sequences converging to pseudoconvex boundary points of finite type in the sense of [10], including uniformly $\Lambda$-tangential and spherically $\frac{1}{2m}$-tangential convergence patterns.
@@ 1. Introduction
We would like to emphasize that the squeezing function $\sigma_{\Omega}$ has a fundamentally global character: its value at a point $p\in\Omega$ depends on holomorphic embeddings of the entire domain $\Omega$ into the ball, not merely on the local geometry near $p$. Therefore, for a boundary point $\xi_{0}\in\partial\Omega$, it is more natural to study the boundary behavior of the localized squeezing function, that is, $\sigma_{\Omega\cap U}$ for a suitable neighborhood $U$ of $\xi_{0}$, rather than that of $\sigma_{\Omega}$ itself. For the localization of the squeezing function, we refer the reader to [16, 29, 39].
In this paper, we investigate the phenomenon of the localized squeezing function approaching $1$ along sequences converging to pseudoconvex boundary points by the scaling technique, introduced by S. Pinchuk (cf. [38]). For a smoothly bounded planar domain $D$ and $p\in\partial D$, one always has $\lim\limits_{z\to p}\sigma_{D}(z)=1$ (see [11]). For pseudoconvex domains of higher dimension, this result holds when $D$ is a bounded strongly pseudoconvex domain in $\mathbb{C}^{n+1}$ (cf. [12]), or when $p$ is globally strongly convex in the sense of [12], or when $p$ is a spherically extreme boundary point in the sense of [24].
The first aim of this paper is to prove the following theorem, which shows that the localized squeezing function approaches $1$ along any sequence converging to a strongly pseudoconvex boundary point.
@@
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n+1}$ and $\xi_{0}\in\partial\Omega$. If $\partial\Omega$ is $\mathcal{C}^{2}$ strongly pseudoconvex at $\xi_{0}$, then we have $\lim\limits_{z\to\xi_{0}}\sigma_{\Omega\cap U_{0}}(z)=1$, where $U_{0}$ is a sufficiently small neighborhood of $\xi_{0}$.
@@
We note that A. Zimmer [45] first proved that if $\lim_{z\to\partial D}\sigma_{D}(z)=1$ for a $\mathcal{C}^{\infty}$-smooth bounded convex domain $D$, then the domain is necessarily strictly pseudoconvex. Subsequently, in contrast to Problem 1, a bounded convex $\mathcal{C}^{2}$-smooth domain $\Omega\subset\mathbb{C}^{n+1}$ that is not strongly pseudoconvex was constructed in [18], satisfying $\lim_{z\to\partial\Omega}\sigma_{\Omega}(z)=1$. Moreover, the authors [35, Example $1.2$] showed that $\sigma_{E_{1,2}}(\eta_{j})\to 1$ as $j\to\infty$ for the sequence
@@
converging tangentially to the non-strictly pseudoconvex boundary point $(0,1)$, where $E_{1,2}:=\{(z_{1},z_{2})\in\mathbb{C}^{2}\colon|z_{2}|^{2}+|z_{1}|^{4}<1\}$ (see also [35, Theorem $1.10$] for the general case). In what follows, we demonstrate that this phenomenon occurs for much larger classes of domains.
The second aim of this paper is to prove the following theorem, which tells us that the localized squeezing function approaches $1$ along a sequence converging uniformly $\Lambda$-tangentially to a strongly $h$-extendible boundary point (cf. Definition 4.2 and Definition 4.1, respectively).
@@
Let $\Omega$ be a bounded pseudoconvex domain in $\mathbb{C}^{n+1}$ with $C^{\infty}$-smooth boundary. Let $\xi_{0}\in\partial\Omega$ be strongly $h$-extendible with Catlin’s finite multitype $(2m_{1},$ $\ldots,2m_{n},1)$ and let $\Lambda=(1/2m_{1},\ldots,1/2m_{n})$ (see Definition 4.2). If $\{\eta_{j}\}\subset\Omega$ is a sequence converging uniformly $\Lambda$-tangentially to $\xi_{0}\in\partial\Omega$ (see Definition 4.1), then $\sigma_{\Omega\cap U_{0}}(\eta_{j})\to 1$ as $j\to\infty$, where $U_{0}$ is a sufficiently small neighborhood of $\xi_{0}$.
@@
Now we turn to bounded pseudoconvex domains in $\mathbb{C}^{2}$. First of all, let us consider the following problem, first posed in [18, Problem $4.3$].
@@
Let $\Omega$ be a bounded domain in $\mathbb{C}^{2}$ and $\xi_{0}\in\partial\Omega$. Suppose that $\partial\Omega$ is $\mathcal{C}^{\infty}$-smooth, pseudoconvex and of D’Angelo finite type near $\xi_{0}$. If $\{\eta_{j}\}\subset\Omega$ is a sequence converging spherically $\frac{1}{2m}$-tangentially to $\xi_{0}\in\partial\Omega$, then $\sigma_{\Omega\cap U_{0}}(\eta_{j})\to 1$ as $j\to\infty$, where $2m$ is the type of $\partial\Omega$ at $\xi_{0}$, where $U_{0}$ is a sufficiently small neighborhood of $\xi_{0}$.
@@ 3.1. Geometry of strongly pseudoconvex hypersurfaces
Shrinking $U_{0}$ if necessary, we may assume that $\Phi_{\eta_{j}^{\prime}}(U_{0}\cap\Omega)$ is contained in the domain $\mathcal{E}:=\big\{(\tilde{z},\tilde{w})\in\mathbb{C}^{n+1}\colon A\,\mathrm{Re}(\tilde{w})+|\tilde{z}|^{2}<B(|\tilde{w}||\tilde{z}|+|\tilde{w}|^{2})\big\})$ for all all sufficiently large $j$, where $A,B$ ($A>1$) are positive constants, independent of $j$.
@@
Therefore, passing to a subsequence if necessary, we may assume that the sequence $\Omega_{j}:=\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(\Omega\cap U_{0})$ converges normally to the Siegel half-space
@@
Next, let us consider the sequence of biholomorphic map $f_{j}:=\Psi\circ\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}\colon\Omega\cap U_{0}\to\Psi(\Omega_{j})$. Then, for any $(z,w)\in U_{0}\cap\Omega$, the point $(z^{\prime},w^{\prime}):=\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(z,w)\in\Omega_{j}$ satisfies
$\mathrm{Re}(w^{\prime})\gtrsim-\dfrac{1}{\epsilon_{j}};\quad|z_{k}^{\prime}|\lesssim\frac{1}{\sqrt{\epsilon_{j}}},\quad 1\leq k\leq n.$
Moreover, for all sufficiently large $j$, we have
$\Omega_{j}\subset\Delta_{j}(\mathcal{E})=\{(^{\prime}z^{\prime},^{\prime}w^{\prime})\in\mathbb{C}^{n+1}\colon A\,\mathrm{Re}(w^{\prime})+|\tilde{z}|^{2}<B(\sqrt{\epsilon_{j}}|w^{\prime}||z^{\prime}|+\epsilon_{j}|w^{\prime}|^{2})\}.$
Consequently, $|z^{\prime}|^{2}\lesssim|\mathrm{Re}(w^{\prime})|$ for all $(z^{\prime},w^{\prime})\in\Omega_{j}$. In addition, for $(\tilde{z},\tilde{w})=\Psi(z^{\prime},w^{\prime})$ with $(z^{\prime},w^{\prime})\in\Omega_{j}$, we have the estimates
$|\tilde{w}+1|=\Big|\dfrac{2}{1-w^{\prime}}\Big|\lesssim\dfrac{1}{|\mathrm{Re}(w^{\prime})|},\quad|\tilde{z}_{k}|=\Big|\frac{2z^{\prime}_{k}}{1-w^{\prime}}\Big|\lesssim\frac{\sqrt{|\mathrm{Re}(w^{\prime})|}}{|\mathrm{Re}(w^{\prime})|-1}\lesssim\frac{1}{\sqrt{|\mathrm{Re}(w^{\prime})|}}$
for $1\leq k\leq n$. Therefore, for any sufficiently small $\epsilon>0$, there exist $M>0$ and $j_{0}\in\mathbb{N}_{\geq 1}$ such that the image under $\Psi$ of the domain
$\Big\{(z^{\prime},w^{\prime})\in\Omega_{j}\colon\mathrm{Re}(w^{\prime})<-M\Big\}$
is contained in $B\big((0^{\prime},-1),\epsilon/2\big)$ for all $j\geq j_{0}$. In addition, the sequence $\big\{(z^{\prime},w^{\prime})\in\overline{\Omega_{j}}\colon\mathrm{Re}(w^{\prime})\geq-M\big\}$ converges to the set $\big\{(z^{\prime},w^{\prime})\in\mathcal{U}_{n+1}\colon\mathrm{Re}(w^{\prime})\geq-M\big\}$ in the Hausdorff sense. Furthermore, we have
$f_{j}(\eta_{j})=\Psi\circ\Theta(0^{\prime},-1)=(0^{\prime},0)\quad\text{for all }j\in\mathbb{N}_{\geq 1}.$
This yields that for any sufficiently small $\epsilon>0$, there exists $j_{0}\in\mathbb{N}_{\geq 1}$ such that
$B\big((0^{\prime},0),1-\epsilon\big)\subset f_{j}(U_{0}\cap\Omega)\subset B\big((0^{\prime},0),1+\epsilon\big),\quad\forall j\geq j_{0},$
and hence
$\sigma_{U_{0}\cap\Omega}(\eta_{j})\geq\frac{1-\epsilon}{1+\epsilon},\quad\forall j\geq j_{0},$
Since $\epsilon>0$ is arbitrary, we conclude that $\lim_{j\to\infty}\sigma_{U_{0}\cap\Omega}(\eta_{j})=1$, and thus the proof of Theorem 1.1 is complete. $\Box$
@@ 4.1. $\Lambda$-tangential convergence
and we consider the sequence $\{Q_{j}\}$ of automorphisms of $\mathbb{C}^{n+1}$, given by
@@
One notes that $M_{H}$ is also the limit of a sequence of the pseudoconvex domains $T_{j}(\Omega\cap U_{0})$. Hence, $M_{H}$ is also pseudoconvex, and thus $H$ is plurisubharmonic. In addition, it follows directly from Lemma 4.1 that $H$ is positive definite. Therefore, there exits a biholomorphism $\Theta\colon M_{H}\to\mathcal{U}_{n+1}$, where $\mathcal{U}_{n+1}$ is the Siegel half-space, given by
@@
Now let us consider the sequence of biholomorphic maps $f_{j}:=\Psi\circ\Theta\circ\Delta_{j}\circ Q_{j}\circ L_{\eta_{j}^{\prime}}\colon\Omega\to f_{j}(\Omega)=\Psi\circ\Theta(\Omega_{j})$. Then the sequences $f_{j}(\Omega\cap U_{0})$ and $f_{j}(\partial\Omega\cap U_{0})$ converge normally to $\mathbb{B}^{n+1}$ and $\partial\mathbb{B}^{n+1}$, respectively. Let us denote $\Omega_{j}:=\Theta\circ\Delta_{j}\circ Q_{j}\circ L_{\eta_{j}^{\prime}}(\Omega\cap U_{0})$ for all $j\in\mathbb{N}_{\geq 1}$. Thus, for any $(z,w)\in U_{0}\cap\Omega$, the point $(z^{\prime},w^{\prime}):=\Theta\circ\Delta_{j}\circ Q_{j}\circ L_{\eta_{j}^{\prime}}(z,w)\in\Omega_{j}$ satisfies
$\mathrm{Re}(w^{\prime})\gtrsim-\dfrac{1}{\epsilon_{j}};\quad|z_{k}^{\prime}|\lesssim\frac{1}{\tau_{jk}},\quad 1\leq k\leq n.$
As in the proof of Theorem 1.1, by [36, Lemma 3.2 and Lemma 3.3] and the equation (3), shrinking $U_{0}$ if necessary we may assume that
$|z^{\prime}|^{2}\lesssim|\mathrm{Re}(w^{\prime})|$
for all $(z^{\prime},w^{\prime})\in\Omega_{j}$, where the constant is independent of $j$. Moreover, for $(\tilde{z},\tilde{w})=\Psi(z^{\prime},w^{\prime})$ with $(z^{\prime},w^{\prime})\in\Omega_{j}$, we have the estimates
$|\tilde{w}+1|=\Big|\dfrac{2}{1-w^{\prime}}\Big|\lesssim\dfrac{1}{|\mathrm{Re}(w^{\prime})|},\quad|\tilde{z}_{k}|=\Big|\frac{2z^{\prime}_{k}}{1-w^{\prime}}\Big|\lesssim\frac{\sqrt{|\mathrm{Re}(w^{\prime})|}}{|\mathrm{Re}(w^{\prime})|-1}\lesssim\frac{1}{\sqrt{|\mathrm{Re}(w^{\prime})|}}$
for $1\leq k\leq n$. Therefore, for any sufficiently small $\epsilon>0$, there exist $M>0$ and $j_{0}\in\mathbb{N}_{\geq 1}$ such that the image under $\Psi$ of the domain
$\Big\{(z^{\prime},w^{\prime})\in\Omega_{j}\colon\mathrm{Re}(w^{\prime})<-M\Big\}$
is contained in $B\big((0^{\prime},-1),\epsilon/2\big)$ for all $j\geq j_{0}$. In addition, the sequence $\big\{(z^{\prime},w^{\prime})\in\Omega_{j}\colon\mathrm{Re}(w^{\prime})\geq-M\big\}$ converges to the set $\big\{(z^{\prime},w^{\prime})\in\mathcal{U}_{n+1}\colon\mathrm{Re}(w^{\prime})\geq-M\big\}$ in the Hausdorff sense. Furthermore, since $\Theta(0^{\prime},-1)=(0^{\prime},-1)$ and $\Psi(0^{\prime},-1)=(0^{\prime},0)$, we have
@@
This yields that for any sufficiently small $\epsilon>0$, there exists $j_{0}\in\mathbb{N}_{\geq 1}$ such that
$B\big((0^{\prime},0),1-\epsilon\big)\subset F_{j}(U_{0}\cap\Omega)\subset B\big((0^{\prime},0),1+\epsilon\big),\quad\forall j\geq j_{0},$
where $F_{j}(\cdot):=f_{j}(\cdot)-f_{j}(\eta_{j})$ for all $j\geq j_{0}$. Since $F_{j}(\eta_{j})=0$, it follows that
$\sigma_{\Omega\cap U_{0}}(\eta_{j})\geq\frac{1-\epsilon}{1+\epsilon},\quad\forall j\geq j_{0}.$
Since $\epsilon>0$ is arbitrary, we conclude that $\lim_{j\to\infty}\sigma_{\Omega\cap U_{0}}(\eta_{j})=1$, and thus the proof of Theorem 1.2 is complete. $\Box$
@@
$\rho(z_{1},z_{2},w)=\mathrm{Re}(w)+|z_{1}|^{4}+|z_{1}|^{2}|z_{2}|^{4}+|z_{2}|^{8}.$
Then we have
$\liminf_{j\to\infty}\sigma_{\Omega\cap U_{0}}(\eta_{j})>0,$
where $\eta_{j}=\big(\frac{1}{j^{1/4}},\frac{1}{j^{3/8}},-\frac{1}{j}-\frac{2}{j^{2}}-\frac{1}{j^{3}}\big)\in\Omega,\;\forall j\in\mathbb{N}_{\geq 1}$, and $U_{0}$ is a sufficiently small neighborhood of $(0,0,0)$.
@@
This yields that the sequence of domains $\Omega_{j}:=\phi_{{\eta}_{j}}(\Omega\cap U_{0})$, where $U_{0}$ is a sufficiently small neighborhood of $(0,0,0)$, converges normally to the following model
@@
Finally, let us consider the sequence of biholomorphic map $f_{j}:=\Psi\circ\Theta\circ\Phi_{\eta_{j}}\colon\Omega\to f_{j}(\Omega)$. Since $\Psi\circ\Theta\circ\phi_{{\eta}_{j}}(\eta_{j}^{\prime})=\Psi(0,1,-1)=(0,1,0)\in\partial\mathcal{D}_{1,1,2}$ and $\Psi\circ\Theta\circ\phi_{{\eta}_{j}}(\eta_{j})=\Psi(0,1,-2)=\Big(0,\sqrt{\dfrac{2}{3}},-\dfrac{1}{2}\Big)\in\mathcal{D}_{1,1,2}$. Furthermore, we have $\Psi(z_{1},z_{2},w)\to(0,0,-1)$ as $\mathcal{E}_{1,1,2}\ni(z_{1},z_{2},w)\to\infty$. As in the proof of Theorem 1.2, we conclude that $f_{j}(\Omega\cap U_{0})$ and $f_{j}(\partial\Omega\cap U_{0})$ converge to $\mathcal{D}_{1,1,2}$ and $\partial\mathcal{D}_{1,1,2}$ in the Hausdorff sense, respectively. Therefore, this implies that
$\sigma_{\Omega\cap U_{0}}(\eta_{j})=\sigma_{f_{j}(\Omega\cap U_{0})}\Big(0,\sqrt{2/3},-1/2\Big)>\dfrac{1}{2}\dfrac{\mathrm{dist}\Big(\big(0,\sqrt{2/3},-1/2\big),\partial\mathcal{D}_{1,1,2}\Big)}{\mathrm{diam}(\mathcal{D}_{1,1,2})}>0$
for any $j$ big enough, where $\mathrm{diam}(\mathcal{D}_{1,1,2})$ denotes the diameter of the domain $\mathcal{D}_{1,1,2}$. Hence, the proof is now complete. ∎
@@
Consider the domain $E_{1,2,4}$ and the sequence $\{\eta_{j}\}\subset E_{1,2,4}$ as in Proposition 4.2. Then we have $\sigma_{E_{1,2,4}}(\eta_{j})\not\to 1$ as $j\to\infty$, contrary to Theorem 1.2. Indeed, suppose, for the sake of contradiction, that $\sigma_{E_{1,2,4}}(\eta_{j})\to 1$ as $j\to\infty$, then by the argument as in the proof of [33, Theorem $2.1$] the unit ball $\mathbb{B}^{n+1}$ is biholomorphically equivalent to $\mathcal{D}_{1,1,2}=\left\{(z_{1},z_{2},w)\in\mathbb{C}^{3}\colon|w|^{2}+|z_{1}|^{2}+|z_{2}|^{4}<1\right\}$. Therefore, we arrive at a contradiction, as $\mathcal{D}_{1,1,2}$ is not homogeneous.